OpenStudy (anonymous):

determine if (-5/6,-1) is a solution to 6x-5y=.10

OpenStudy (lalaly):

substitute 6(-5/6) - 5(-1)=10 -5+5=0 its not a solution

OpenStudy (anonymous):

(-5/6,-1) is the (x,y) values. Subsitute x=-5/6 and y=-1 into 6x-5y=10. So 6(-5/6) - 5(-1) ----> this should give 10 6(-5/6) - 5(-1)=0 ≠10 Hence (-5/6,-1) is not a solution to 6x-5y=10

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

No.\[6\left(-\frac{5}{6}\right)-5(-1)=0 \]